DIRC Operations Weekly Report

Period: Wednesday 2007/5/30 - Wednesday 2007/6/6


Oncall: Jose Benitez
Fast Monitoring: Jose Benitez, Alejandro Perez

Contents


    PEP Activities     BaBar Activities     DIRC activities     Backgrounds     Scheduled activities     Pending Work     Links

Luminosities this week

PEP peak lumi: ~9.5 x 1033 (last week: ~8.5 x 1033)

Run 6 PEP delivery: 4.8 fb-1 (last weeks: 2.88 fb-1)

BaBar integrated lumi:4.6 fb-1 (last weeks: 2.81 fb-1)

Run 6 total delivery: ~42.1 fb-1

Run 6 total BaBar recorded: ~40.5 fb-1 including 0 fb-1 off-peak data

PEP activities/issues this week

  • Several periods of high backgrounds, some associated with vacuum problems.
  • Generally delivered decent luminosity but with high backgrounds

    BaBar activities/issues this week

  • Several periods of high dead time due to bad backgrounds.

    DIRC activities/issues this week

  • 6/31: Got paged due to persistent high backgrounds in the DIRC.
  • several noisy channels coming and going in sectors 0 and 6.